OpenStudy (anonymous):

A toy train is pushed forward and released at xi = 1.0 m with a speed of 1.0 m/s. It rolls at a steady speed for 2.5 s, then one wheel begins to stick. The train comes to a stop 6.0 m from where it was released. What is the train's acceleration after its wheel begins to stick?

OpenStudy (anonymous):

So the train is moving initially at 1.0 m/s If it rolls for 2.5 s at this speed, the product gives us the distance: x = v * t = (1.0) * (2.5) = 2.5 m The train moves 2.5 m before it's wheel gets stuck. This means that there is 6 - 2.5 m or 3.5 m left until the train stops. So we need to find how much the train decelerates over 3.5 m to go from 1.0 m/s to 0.0 m/s. kinematic eq: Vf^2 = Vi^2 + 2a(Xf - Xi) Vf = 0 m/s Vi = 1.0 m/s a = ? m/s^2 <-- we need to find this Xf = 3.5 m Xi = 0 m <-- we'll take the point where it starts to get stuck as the new 0 position Now we solve: 0 = 1^2 + (2)(a)(3.5 - 0) 0 = 1 + 7a -1 = 7a a = -1/7 = -0.14 m/s^2 The negative sign indicates the train is slowing down, which is correct! The train accelerates at -0.14 m/s^2 after it's wheel starts to stick.

OpenStudy (anonymous):

If it is easier to understand, you could replace Xf and Xi with what you actually have: Xf = 6 + Xi = 6 + 1 = 7 m new Xi = old Xi + 2.5 = 1 + 2.5 = 3.5 m Notice, 7 - 3.5 is the same as what we did above (3.5 - 0).